### v2

# COLT: Enhancing Video Large Language Models with Continual Tool Usage

Abstract
The success of Large Language Models (LLMs) has significantly propelled the research of video understanding. To harvest the benefits of well-trained expert models (i.e., tool), video LLMs prioritize the exploration of tool usage capabilities. Existing methods either prompt closed-source LLMs or employ the instruction tuning paradigm for tool-use finetuning. These methods, however, assume an established repository of fixed tools and struggle to generalize to real-world environments where tool data is perpetually evolving and streaming in. To this end, we propose to enhance open-source video LLMs with COntinuaL Tool usage (termed COLT), which automatically acquires tool-use ability in a successive tool stream without suffering “catastrophic forgetting” of the past learned tools. Specifically, our COLT incorporates a learnable tool codebook as a tool-specific memory system. Then relevant tools are dynamically selected based on the similarity between user instruction and tool features within the codebook. To unleash the tool usage potential of video LLMs, we collect a video-centric tool-use instruction tuning dataset VideoToolBench. Extensive experiments on both previous video LLM benchmarks and the tool-use-specific VideoToolBench dataset demonstrate the state-of-the-art performance of our proposed COLT.
Introduction
Large Language Models (LLMs), e.g., GPT (Achiam et al. 2023; Brown et al. 2020), PaLM (Anil et al. 2023) and LLaMA (Touvron et al. 2023a, b), have exhibited remarkable success in understanding user instructions, aligning with human intents and generating trustworthy responses. Taking one step further, recent endeavors in multi-modal LLMs (Li et al. 2023a; Alayrac et al. 2022; Zhu et al. 2023b; Liu et al. 2024a; Zhang et al. 2023a; Wu et al. 2023; Li et al. 2024b; Zhan et al. 2024) augment LLMs with the capacity of multi-modal perception, which has aroused significant research interest within the academic community and beyond.
Figure 1: Our proposed COLT continually learns to invoke tools from tool-stream data without catastrophic forgetting. Benefiting from tool usage, COLT (a) is adept at dynamic content understanding and (b) supports flexible generation compared to existing methods (Lin et al. 2023a; Li et al. 2023c). The incorrect parts of responses are marked in red.
Figure 2: (a) Agent-based methods bootstrap closed-source LLMs via delicately designed prompts; (b) Instruction tuning with fixed tool-use dataset; (c) Instruction tuning with steam tool-use dataset (Ours); (d) Average tool calling accuracy on VideoToolBench vs. # learned tools. Sequential training denotes training on a sequence of tasks independently, one after the other.
Despite the progress in image-based LLMs, advancing LLMs’ capacity to comprehend video data (Lin et al. 2023a; Zhang, Li, and Bing 2023; Li et al. 2023b; Maaz et al. 2023; Li, Wang, and Jia 2023; Jin et al. 2024; Munasinghe et al. 2023; Liu et al. 2024b; Lin et al. 2023c; Jin et al. 2023; Liu et al. 2023b) still remains a demanding pursuit. The overwhelming majority of video LLM methods follow the instruction-tuning paradigm. Building upon open-source LLMs (Chiang et al. 2023; Touvron et al. 2023a, b), they entail end-to-end training on the instruction-tuning dataset generated by GPT (Achiam et al. 2023; Brown et al. 2020). These methods, however, demonstrate limited ability to comprehend video actions and present responses with hallucinations (e.g., non-existent “keyboard click” actions). Besides, most of them lack the capability of any-to-any generation, e.g., video object segmentation in Figure 1 (b).
To alleviate this, agent-based methods (Gao et al. 2023; Qin et al. 2023; Yang et al. 2024b; Fan et al. 2024; Wang et al. 2024b) are introduced to advocate the tool-use capabilities of LLMs. As shown in Figure 2 (a), they mostly bootstrap closed-source LLMs (Achiam et al. 2023; Brown et al. 2020) to decompose the user instructions into more manageable sub-tasks and incorporate external tools to assist. These training-free methods require delicately designed prompts and lengthy context windows for instructions (Liu et al. 2023c). Instead, recent attempts (Liu et al. 2023c; Qin et al. 2023) explore developing tool-specific instruction tuning data and elicit the tool-use capabilities of open-source LLMs (c.f. Figure 2 (b)). However, these methods typically assume a pre-set repository containing fixed tools, which is difficult to generalize into the real-world scenario where tool data arrives in a never-ending stream. In such case, the conventional sequential training, i.e., training on new data with the pre-trained weights on the previous tool data as initialization, leads to significant catastrophic forgetting. of prior tool-usage knowledge. For example in Figure 2 (d), the tool-call accuracy of sequential training rapidly decreases as the number of tools learned increases.
In this paper, we propose to enhance open-source video LLMs with COntinuaL Tool usage (dubbed as COLT). By “continual”, we mean that video LLMs possess the inherent capacity to learn and automatically activate tool functionality in a successive tool-stream format. To achieve this, two critical issues arise: 1) how to build a tool-specific memory to mitigate catastrophic forgetting of past knowledge? One intuitive idea is to store a few past tool data and replay them with the new tool data for continual training. However, the requisite memory expands proportionally with the increase in the number of tools and previous data may be unavailable due to privacy constraints. Instead, we set up a tool codebook consisting of learnable prompts to store tool-specific information in a more concise manner; 2) how to exploit relevant tools based on input user instructions? For tool activation, we compute the cosine similarity between the embedded user instructions and each prompt within the tool codebook. The highest-response prompts are then selected to “instruct” the model to invoke appropriate tools.
Although several video LLM benchmarks (Maaz et al. 2023; Li et al. 2023c) have been proposed, the community still lacks the benchmark for video-centric tool-use instruction tuning. To this end, we collect VideoToolBench to facilitate tool-use capabilities for open-source video LLMs. Specifically, we collect a set of video specialist models and prompt GPT (Achiam et al. 2023) to generate diverse instructions for tool calls. Then the tool execution results are used to form the instruction-following dataset. Due to our continual learning setting, the tool list can be easily extended, making our COLT a lifelong learning video LLM.
In summary, our contributions are in three-folds:
  • We proposed COLT, a video LLM with continual tool usage. By maintaining a tool codebook, COLT incrementally learns new tools without catastrophic forgetting.
  • A tool-using instruction-following dataset VideoToolBench is introduced to unlock the potential for tool usage within video LLMs.
  • Experiments on both existing video LLM benchmarks and VideoToolBench have manifested the state-of-the-art performance of COLT. For example on MSRVTT-QA (Chen and Dolan 2011), our COLT boosts the previous SOTA method (Li et al. 2023c) by 8.2% on the accuracy of zero-shot video-question answering.
Related Work
Video LLMs. Recent successes of LLMs (Achiam et al. 2023; Anil et al. 2023) have shed light on the burgeoning proliferation of video LLMs (Lin et al. 2023a; Maaz et al. 2023; Gao et al. 2023). The mainstream video LLMs follow the instruction-tuning paradigm. Building upon open-source LLMs (Chiang et al. 2023; Touvron et al. 2023a, b), this kind of methods adapts the pre-trained video features into LLM understandable representations via multi-layer perception projector (Lin et al. 2023a; Maaz et al. 2023; Li, Wang, and Jia 2023; Munasinghe et al. 2023; Liu et al. 2024b; Jin et al. 2023; Liu et al. 2023b), Q-former (Zhang, Li, and Bing 2023; Li et al. 2023b), or discretization tokenizer (Jin et al. 2024). These methods, however, fail to generalize to broader video understanding tasks which may require flexible input and output formats (Wu et al. 2023; Zhan et al. 2024). Several methods attempt to achieve this by employing additional functional modules (Jin et al. 2024; Munasinghe et al. 2023) (e.g., grounding (Munasinghe et al. 2023) or diffusion modules (Jin et al. 2024)), which is not flexible enough to adapt to diverse video understanding needs. To this end, another stream of agent-based methods (Gao et al. 2023; Qin et al. 2023; Yang et al. 2024b; Fan et al. 2024; Wang et al. 2024b) is proposed, where multi-modal agents mostly bootstrap closed-source LLM (Achiam et al. 2023; Brown et al. 2020) to decompose solution paths and call external tools. This kind of method relies heavily on delicately designed prompts and may fail to acquire the usage of a novel tool. Therefore, how to implement automatic invocation of related tools in video LLMs under a non-stationary tool stream remains to be solved. This paper combines the strengths of the above two methodologies to empower video LLMs with automatic and continual tool-use ability.
Continual Learning. Continual learning (Wang et al. 2024a; Lee et al. 2017; McCloskey and Cohen 1989) refers to the ability to incrementally acquire, update, and accumulate knowledge throughout the model lifetime without catastrophically forgetting previously learned information. Conceptually, four varieties of methodologies are posited. Regularization-based approaches (Kirkpatrick et al. 2017; Li and Hoiem 2017; Feng, Wang, and Yuan 2022; Yang et al. 2024a) strike the balance between the old and new tasks by adding explicit regularization terms. Architecture-based approaches  (Yoon et al. 2018; Li et al. 2019; Ke, Liu, and Huang 2020) isolate model parameters for different tasks. Rehearsal-based methods (Bonicelli et al. 2022; Chen and Chang 2023; Lin et al. 2023b) typically use a memory buffer to store several training samples from previous classes, which are used to approximate and recover old data distributions. Prompt-based methods (Wang et al. 2022c; Smith et al. 2023; Wang et al. 2022b; Li et al. 2024a) usually construct task-adaptive prompts and select appropriate prompts during inference. This kind of method is rehearsal-free and thus more computationally efficient. L2P (Wang et al. 2022c) introduces the concept of prompt pool and select prompts by a query-key mechanism. To overcome the separate optimization issue of L2P (Wang et al. 2022c), CODA-Prompt (Smith et al. 2023) assembles learnable prompts with input-conditioned weights. DualPrompt (Wang et al. 2022b) and KC-Prompt (Li et al. 2024a) set up prompt pools to respectively encode task-invariant and tasks-specific knowledge.
Borrowing the favorable rehearsal-free merit, our COLT inventively devises a video LLM with the continual tool-use learning capability. We differ from current prompt-based methods (Wang et al. 2022c; Smith et al. 2023; Wang et al. 2022b; Li et al. 2024a) in the following two aspects. Firstly, our focus is on the more complex task of multi-modal language generation as opposed to the basic image classification task. Secondly, our COLT employs straight-through (Van Den Oord, Vinyals et al. 2017; Bengio, Léonard, and Courville 2013) gradient estimation to mitigate the optimization challenges encountered in previous prompt-based approaches (Wang et al. 2022c; Smith et al. 2023; Wang et al. 2022b; Li et al. 2024a).
Dataset Construction
Dataset Structure. Each instance in VideoToolBench contains two rounds of conversations between human and gpt. The first round includes human instructions related to the video content and the LLM responses on choosing appropriate tools; The second round contains the execution results of the corresponding tools and the final responses from LLMs. Specifically, we follow (Yao et al. 2022; Yang et al. 2023; Liu et al. 2023c) to unify the response format of GPT into three fields including thought, action and value to mimic human-like task-solving procedure.
Table 1: Tool repository of the proposed VideoToolBench including single and compositional tools.
Tool Specialist Model
single tool
Action Recognition (AR) VideoMAE (Tong et al. 2022)
Dense Video Caption (DVC) PDVC (Wang et al. 2021)
Temporal Action Localization (TAL) InternVideo (Wang et al. 2022a)
Optical Character Recognition (OCR) EasyOCR (JaidedAI 2023)
Automatic Speech Recognition (ASR) Whisper (Radford et al. 2023)
Video Relation Detection (VRD) VidVRD (Shang et al. 2017)
Video Object Segmentation (VOS) VisTR (Wang et al. 2021)
Text-to-Video Generation (T2V) T2V (Khachatryan et al. 2023)
multi.
AR + ASR VideoMAE + Whisper
AR + VOS VideoMAE + VisTR
Dataset Construction. This dataset is constructed using GPT-3.5-turbo (Achiam et al. 2023) with self-instruction. The involved tools [^1] In this paper, we use tool to represent the general skills and specialist model to denote the specific model. incorporate both video understanding and generation tasks. The prompt and in-context learning cases are available in supplementary materials. As shown in Table 1, we collect ten tools including eight single tool and two compositional tool. The tool list can be easily extended using a similar dataset construction manner. We initially generate 5,000 instruction-following samples for each tool. We conduct both data format checks and manual verification of semantic meanings to filter out error data. After a thorough examination, the dataset is partitioned into distinct training and testing splits, adhering to a proportionality ratio of 9:1. The full statistics of the final VideoToolBench are summarized in supplementary materials.
Method
Our COLT learns the continual tool usage from the stream instruction-tuning dataset $\{\mathcal{D}^{t}\}_{t=1}^{T}$, each $\mathcal{D}^{t}$ containing triplets of the visual data $\mathbf{X}_{\mathbf{v}}^{t}$, user instructions $\mathbf{X}_{\mathbf{w}}^{t}$ and the response sequence $\mathbf{X}_{\mathbf{a}}^{t}$. For clarity, we elaborate on the architecture and training strategy for one single dataset $\mathcal{D}^{t}$ and the superscripts of $\mathbf{X}_{\mathbf{v}}^{t}$, $\mathbf{X}_{\mathbf{w}}^{t}$ and $\mathbf{X}_{\mathbf{a}}^{t}$ are omitted.
Figure 3: An overview of COLT. Stage 1 aligns the visual and textual modalities through the individual training of the linear projector ${f}_{\boldsymbol{\delta}}$; In stage 2 and stage 3, the prompt within tool codebook $\mathbf{P}$ is adaptively selected according to the cosine similarity with the query feature $\mathbf{H}_{\mathbf{q}}$.
Architecture
Visual & Textual Embedding. Following (Lin et al. 2023a), we adopt LanguageBind (Zhu et al. 2023a) vision encoder ${g}(\cdot)$ to extract visual features. Then a linear projection layer ${f}_{\boldsymbol{\delta}}(\cdot)$ parameterized by $\boldsymbol{\delta}$ is applied to compress the visual information into LLM understandable feature space.
$\mathbf{H}_{\mathbf{v}}={f}_{\boldsymbol{\delta}}({g}(\mathbf{X}_{\mathbf{v}})),$ (1)
where $\mathbf{H}_{\mathbf{v}}\in\mathbb{R}^{P\times C}$ denotes the visual embeddings, with $P$ and $C$ respectively representing the patch number and the feature dimension. For input user instructions $\mathbf{X}_{\mathbf{w}}$, we adopt the widely used BPE tokenizer (Sennrich, Haddow, and Birch 2016) to obtain the textual embeddings $\mathbf{H}_{\mathbf{w}}\in\mathbb{R}^{S\times C}$, where $S$ is the textual token number.
Tool Codebook & Query Encoder. As shown in Figure 3, we set up a tool codebook $\mathbf{P}$ as the tool-specific memory for continual tool usage. The codebook consists of NN learnable tool prompts, i.e., $\mathbf{P}=\{\mathbf{P}_{n}\}_{n=1}^{N}$ with $\mathbf{P}_{n}\in\mathbb{R}^{C\times 1}$. For each input user instruction $\mathbf{X}_{\mathbf{q}}$, we aim to retrieve relevant tool prompts from the codebook. To achieve this, a query encoder ${f}_{\boldsymbol{\beta}}$ with parameters $\boldsymbol{\beta}$ is firstly employed to encode user instructions $\mathbf{X}_{\mathbf{q}}$ into $\mathbf{H}_{\mathbf{q}}\in\mathbb{R}^{C\times 1}$. Then, we compute the cosine similarities between the query embedding $\mathbf{H}_{\mathbf{q}}$ and each tool prompt $\mathbf{P}_{n}$, $n\in\{1,2,\cdots,N\}$. The tool prompts with top-$K$ highest similarity scores are selected.
$n^{*}={\arg\operatorname{topk}}_{n\in[1,N]}\;(\mathbf{P}_{n}^{\top}\cdot\mathbf{H}_{\mathbf{q}}),$ (2)
where $n^{*}$ is the selected tool prompt index. Then each $\mathbf{P}_{n^{*}}$ is concatenated with the video feature $\mathbf{H}_{\mathbf{v}}$ and user instruction features $\mathbf{H}_{\mathbf{w}}$ and fed into LLMs. The hint contained in the codebook helps alleviate the catastrophic forgetting when facing the tool stream data.
Training
The proposed COLT adopts a three-stage training methodology, with each stage featuring unique trainable parameters $\boldsymbol{\theta}$ and loss functions.
Stage 1: Video-to-text Alignment. This stage aims to train the projection layer ${f}_{\boldsymbol{\delta}}$ parameterized by $\boldsymbol{\delta}$, which acts as a visual tokenizer to align visual signals with pre-trained LLM word embedding. As shown in Figure 3, we freeze all the weights except the projection layer in this stage, i.e., trainable parameters $\boldsymbol{\theta}=\boldsymbol{\delta}$. Given the predicted response sequence $\mathbf{X}_{\mathbf{a}}=\{\mathbf{X}_{\mathbf{a}}^{1},\mathbf{X}_{\mathbf{a}}^{2},\cdots\mathbf{X}_{\mathbf{a}}^{L}\}$ of length $L$, we use the vanilla auto-regressive language modeling (LM) loss to supervise this stage of training:
$\mathcal{L}=\mathcal{L}_{\text{LM}}=-\frac{1}{L}\sum_{i=1}^{L}\;\log p_{\boldsymbol{\theta}}(\mathbf{X}_{\mathbf{a}}^{i}\mid\mathbf{H}_{\mathbf{v}},\mathbf{H}_{\mathbf{w}},\mathbf{X}_{\mathbf{a}}^{<i}),$ (3)
where $\mathbf{X}_{\mathbf{a}}^{<i}$ is the response tokens before the $i$-th token.
Stage 2: Tool Codebook Pre-training. In this stage, we pre-train the tool codebook $\mathbf{P}$ and the query encoder ${f}_{\boldsymbol{\beta}}$ while keeping the other parts frozen, i.e., the trainable parameters are $\boldsymbol{\theta}=\{\mathbf{P},\boldsymbol{\beta}\}$. We firstly select the appropriate tool prompt $\mathbf{P}_{n^{*}}$ via Eq.(2). Then $\mathbf{P}_{n^{*}}$ serves as an additional condition for language modeling as follows:
$\mathcal{L}_{\text{LM}}^{\prime}=-\frac{1}{L}\sum_{i=1}^{L}\;\log p_{\boldsymbol{\theta}}(\mathbf{X}_{\mathbf{a}}^{i}\mid\mathbf{H}_{\mathbf{v}},\mathbf{H}_{\mathbf{w}},{\color[rgb]{0,0.6015625,0.3359375}\mathbf{P}_{n^{*}}},\mathbf{X}_{\mathbf{a}}^{<i}),$ (4)
where $\mathbf{X}_{\mathbf{a}}^{i}$ and $\mathbf{X}_{\mathbf{a}}^{<i}$ is as defined in Eq. (3). However, directly optimizing $\mathcal{L}_{\text{LM}}^{\prime}$ will truncate the gradients w.r.t the query encoder ${f}_{\boldsymbol{\beta}}$ due to the non-derivable $\arg\operatorname{topk}$ operation in Eq. (2). Therefore, we resort to the straight-through estimator (Van Den Oord, Vinyals et al. 2017; Bengio, Léonard, and Courville 2013) to approximate the gradient computation and define the overall loss function as follows:
$\mathcal{L}=\mathcal{L}_{\text{LM}}^{\prime}+\lambda_{1}\underbrace{\left\|\operatorname{sg}\left[\mathbf{H}_{\mathbf{w}}\right]-\mathbf{P}_{n^{*}}\right\|_{2}^{2}}_{\mathcal{L}_{q}}+\lambda_{2}\underbrace{\left\|\mathbf{H}_{\mathbf{w}}-\operatorname{sg}[\mathbf{P}_{n^{*}}]\right\|_{2}^{2}}_{\mathcal{L}_{c}},$ (5)
where $\operatorname{sg}[\cdot]$ stands for the stop-gradient operator that acts as an identity in the forward process and has zero partial derivatives during backward propagation. $\mathcal{L}_{q}$ is the quantisation loss for codebook update by forcing $\mathbf{P}$ towards the user instruction embeddings $\mathbf{H}_{\mathbf{w}}$, $\mathcal{L}_{c}$ is the commitment loss to prevent unrestricted update of codebook embeddings, $\lambda_{1}$ and $\lambda_{2}$ are the balancing weights.
Stage 3: End-to-end Fine-tuning. We keep the vision encoder ${g}(\cdot)$ frozen and finetune remaining parts including the projection layer ${f}_{\boldsymbol{\delta}}$, tool codebook $\mathbf{P}$, textual encoder ${f}_{\boldsymbol{\beta}}$ and LLM parameterized by $\boldsymbol{\phi}$, i.e., the trainable parameters are $\boldsymbol{\theta}=\{\boldsymbol{\delta},\mathbf{P},\boldsymbol{\beta},\boldsymbol{\phi}\}$. The training loss is the same as Eq. (5).
Experiments
Experimental Settings
Training Datasets. In the first stage, we follow Video LLaVA (Lin et al. 2023a) to use LAION-$C$-SBU image subset (Schuhmann et al. 2021) and the filtered CC3M video dataset (Changpinyo et al. 2021) for video-to-text alignment. For the second and third stage, we use the combined dataset of 665K image-text instruction data from LLaVA v1.5 (Liu et al. 2023a) and 100K video-text instruction data from Video-ChatGPT (Maaz et al. 2023). The tool-use instruction data of our collected VideoToolBench is introduced into continual training sequentially. To ensure conformity in data formatting, we reformat all the instruction tuning datasets to the thought-action-value pattern. Refer to supplementary materials for more details.
Table 2: Comparisons with state-of-the-art methods on zero-shot video-question answering. “Acc.” denotes accuracy (%) and “Score” denotes the relative score from 0 to 5 assigned by GPT (Brown et al. 2020). The best performance is in bold and the second best is underlined. The backend LLMs include Vicuna-7B (Chiang et al. 2023) and LLaMA-7B (Touvron et al. 2023a).
Method LLM MSVD-QA MSRVTT-QA ActivityNet-QA
Acc. Score Acc. Score Acc. Score
Video-LLaMA (Zhang, Li, and Bing 2023) Vicuna-7B 51.6 2.5 29.6 1.8 12.4 1.1
VideoChat (Li et al. 2023b) Vicuna-7B 56.3 2.8 45.0 2.5 26.5 2.2
Video-ChatGPT (Maaz et al. 2023) Vicuna-7B 64.9 3.3 49.3 2.8 35.2 2.7
BT-Adapter (Liu et al. 2023b) Vicuna-7B 67.5 3.7 57.0 3.2 45.7 3.2
LLaMA-VID (Li, Wang, and Jia 2023) Vicuna-7B 69.7 3.7 57.7 3.2 47.4 3.3
LLaMA-VID (Li, Wang, and Jia 2023) Vicuna-13B 70.0 3.7 58.9 3.3 47.5 3.3
Video-LLaVA (Lin et al. 2023a) Vicuna-7B 70.7 3.9 59.2 3.5 45.3 3.3
Chat-UniVi (Jin et al. 2023) Vicuna-7B 65.0 3.6 54.6 3.1 45.8 3.2
LLaMA-Adapter (Zhang et al. 2023b) LLaMA-7B 54.9 3.1 43.8 2.7 34.2 2.7
VideoChat2 (Li et al. 2023c) Vicuna-7B 70.0 3.9 54.1 3.3 49.1 3.3
ST-LLM (Liu et al. 2024b) Vicuna-7B 74.6 3.9 63.2 3.4 50.9 3.3
COLT $_{\text{joint}}$ (Ours) Vicuna-7B 78.2 4.2 65.1 3.6 54.7 3.8
COLT $_{\text{5$\times$2}}$ (Ours) Vicuna-7B 75.5 3.9 63.9 3.5 52.5 3.5
COLT $_{\text{10$\times$1}}$ (Ours) Vicuna-7B 74.7 3.9 63.0 3.4 51.2 3.4
Implementation Details. For vision encoder $\boldsymbol{g}(\cdot)$, we choose pre-trained LanguageBind (Zhu et al. 2023a) with ViT-L/14 (Dosovitskiy et al. 2020). The text tokenizer is derived from LLaMA (Touvron et al. 2023a) with a vocabulary size of 32,000 and Vicuna-7B v1.5 (Chiang et al. 2023) is employed as the large language model. We uniformly sample 8 frames from each video and each frame is resized to 224 $\times$ 224. We set the batch size to 256 for the first stage and 128 for the second and third stage. AdamW optimizer is used with a cosine decay schedule. We set learning rate to $\times 10^{-4}, 1×1$$\times 10^{-4}$, and $1\times 10^{-5}$ for three stage training, respectively. The balancing weight $\lambda_{1}$ and $\lambda_{2}$ in Eq. (5) are set to 1 and 0.25. The codebook size NN and selected prompt number $K$ is set to 50 and 3, respectively.
Comparisons with Video LLMs
Table 3: Comparisons (%) with state-of-the-art methods on MVBench. The tasks include Action Sequence (AS), Action Prediction (AP), Action Antonym (AA), Fine-grained Action (FA), Unexpected Action (UA), Object Existence (OE), Object Interaction (OI), Object Shuffle (OS), Moving Direction (MD), Action Localization (AL), Scene Transition (ST), Action Count (AC), Moving Count (MC), Moving Attribute (MA), State Change (SC), Fine-grained Pose (FP), Character Order (CO), Egocentric Navigation (EN), Episodic Reasoning (ER), Counterfactual Inference (CI), and the average of all 20 metrics (AVG). The best performance is in bold and the second best is underlined.
Model AVG AS AP AA FA UA OE OI OS MD AL
InstructBLIP (Dai et al. 2024) 32.5 20.0 16.5 46.0 24.5 46.0 51.0 26.0 37.5 22.0 23.0
LLaVA (Liu et al. 2024a) 36.0 28.0 39.5 63.0 30.5 39.0 53.0 41.0 41.5 23.0 20.5
VideoChatGPT (Maaz et al. 2023) 32.7 23.5 26.0 62.0 22.5 26.5 54.0 28.0 40.0 23.0 20.0
VideoLLaMA (Zhang, Li, and Bing 2023) 34.1 27.5 25.5 51.0 29.0 39.0 48.0 40.5 38.0 22.5 22.5
VideoChat (Li et al. 2023b) 35.5 33.5 26.5 56.0 33.5 40.5 53.0 40.5 30.0 25.5 27.0
VideoChat2text (Li et al. 2023c) 34.7 24.5 27.0 49.5 27.0 38.0 53.0 28.0 40.0 25.5 27.0
VideoChat2 (Li et al. 2023c) 51.1 66.0 47.5 83.5 49.5 60.0 58.0 71.5 42.5 23.0 23.0
GPT-4V (OpenAI 2023) 43.5 55.5 63.5 72.0 46.5 73.5 18.5 59.0 29.5 12.0 40.5
COLT $_{\text{joint}}$ (Ours) 53.4 64.0 65.0 77.0 44.5 54.5 75.5 70.0 34.0 36.5 33.0
COLT $_{\text{5$\times$2}}$ (Ours) 51.8 62.0 61.0 75.0 42.5 52.5 74.5 69.0 33.5 35.0 32.0
COLT $_{\text{10$\times$1}}$ (Ours) 50.6 60.0 60.5 73.5 42.0 51.5 73.5 68.5 32.5 33.0 30.0
Model AVG ST AC MC MA SC FP CO EN ER CI
InstructBLIP (Dai et al. 2024) 32.5 46.5 42.5 26.5 40.5 32.0 25.5 30.0 25.5 30.5 38.0
LLaVA (Liu et al. 2024a) 36.0 45.0 34.0 20.5 38.5 47.0 25.0 36.0 27.0 26.5 42.0
VideoChatGPT (Maaz et al. 2023) 32.7 31.0 30.5 25.5 39.5 48.5 29.0 33.0 29.5 26.0 35.5
VideoLLaMA (Zhang, Li, and Bing 2023) 34.1 43.0 34.0 22.5 32.5 45.5 32.5 40.0 30.0 21.0 37.0
VideoChat (Li et al. 2023b) 35.5 48.5 35.0 20.5 42.5 46.0 26.5 41.0 23.5 23.5 36.0
VideoChat2text (Li et al. 2023c) 34.7 38.5 41.5 27.5 32.5 46.5 26.5 36.0 33.0 32.0 40.0
VideoChat2 (Li et al. 2023c) 51.1 88.5 39.0 42.0 58.5 44.0 49.0 36.5 35.0 40.5 65.5
GPT-4V (OpenAI 2023) 43.5 83.5 39.0 12.0 22.5 45.0 47.5 52.0 31.0 59.0 11.0
COLT $_{\text{joint}}$ (Ours) 53.4 88.0 37.5 53.5 75.5 35.5 42.5 48.0 34.5 42.0 57.5
COLT $_{\text{5$\times$2}}$ (Ours) 51.8 87.0 37.0 52.5 72.0 33.0 40.5 48.0 33.5 40.5 55.0
COLT $_{\text{10$\times$1}}$ (Ours) 50.6 86.0 36.0 51.5 71.5 31.5 39.5 47.5 32.0 38.0 54.0
Table 4: Ablations studies on MVBench. (a) training strategies; (b) training losses; (c) prompt selection mechanisms: T-based and V-based denote tool prompt selection based on text and visual features, respectively; (d) prompt positions.
Stage 2 Stage 3 AVG
50.4
49.2
51.8
$\mathcal{L}_{q}$ $\mathcal{L}_{c}$ AVG
37.7
40.3
51.8
T-based V-based AVG
45.3
51.8
46.4
Position AVG
tool-vision-text 51.7
vision-tool-text 51.7
vision-text-tool 51.8
Table 5: Comparisons with SOTA methods on test split of VideoToolBench. AR, VRD, and ASR denote action recognition, video relation detection and automatic speech recognition, respectively. ”Acc.” denotes accuracy (%) and ”Score” denotes the GPT-evaluated score.
Method AR VRD ASR
Acc. Score Acc. Score Acc. Score
VideoLLaVA 47.7 2.89 6.0 0.82 1.83 0.77
VideoChat2 51.4 3.13 14.0 1.36 3.60 0.87
VideoChat 29.0 2.31 4.5 0.90 2.80 1.28
COLT $_{\text{joint}}$ 77.9 3.84 24.0 1.75 24.2 1.95
COLT $_{\text{5$\times$2}}$ 74.8 3.60 21.3 1.48 22.6 1.50
COLT $_{\text{10$\times$1}}$ 73.5 3.48 19.6 1.27 21.9 1.48
[…]
Evaluation Benchmarks. Our experiments are carried out on both established video LLM benchmarks and self-built tool-using datasets: 1) zero-shot video-question answering. We experiment on commonly used open-ended question-answer datasets: MSVD-QA (Chen and Dolan 2011), MSRVTT-QA (Xu et al. 2016), and ActivityNet-QA (Yu et al. 2019). Following (Li et al. 2023b; Lin et al. 2023a), we use GPT-assisted evaluation to assess the model’s capabilities by reporting the accuracy and relative score; 2) MVBench (Li et al. 2023c). This benchmark consists of 20 demanding video tasks, each comprising 200 samples presented as multiple-choice questions. These tasks offer a thorough and unbiased evaluation of a model’s capacity to comprehend videos. We report the choice accuracy as the metric; 3) VideoToolBench. We build this dataset to probe the abilities enabled by tool proficiency. Since most existing video LLMs only support textual outputs, we select three tools (i.e., action recognition, video relation detection, and automatic speech recognition) and compare our COLT to state-of-the-art video LLMs (Lin et al. 2023a; Li et al. 2023c, b). The GPT-evaluated accuracy and scores are reported.
Performance Analysis. The comparison results on zero-shot video-question answering, MVBench, and VideoToolBench are summarized in Table 2, Table 3, and Table 5, respectively. We can conclude with the following findings. 1) Both COLT $_{\text{joint}}$ and COLT $_{\text{5$\times$2}}$ demonstrate superior performance compared to preceding state-of-the-art methodologies. For example on MSRVTT-QA (c.f. Table 2), COLT $_{\text{joint}}$ remarkably surpasses the previous best performing method VideoChat2 (Li et al. 2023c) by 8.2% on the metric of accuracy; 2) On the test set of VideoToolBench (c.f. Table 5), our COLT significantly outperforms previous methods by a remarkable margin, which demonstrates that our method is more effective in scenarios that require invoking tools to complete; 3) Even with a rather lengthy learning curve, COLT $_{\text{10$\times$1}}$ still achieves comparable performance with previous SOTA methods; 4) The performance of COLT $_{\text{10$\times$1}}$ is slightly worse than COLT $_{\text{5$\times$2}}$, which is consistent with the intuition that tool learning becomes increasingly challenging when facing a lengthening tool-chain.
Comparisons with Continual Learning Methods
Table 6: Comparisons (%) with continual learning methods VideoToolBench. AA and AF denote average accuracy and average forgetting, respectively.
Method Five Groups Ten Groups
AA↑ AF↓ AA↑ AF↓
Sequential 48.6 35.3 42.3 39.7
Rehearsal (10/Tool) 52.9 32.1 48.2 36.3
Rehearsal (30/Tool) 55.2 30.2 49.3 35.5
Rehearsal (50/Tool) 59.3 27.4 50.2 32.7
L2P (Wang et al. 2022c) 72.4 8.3 65.7 12.4
Dual (Wang et al. 2022b) 75.1 6.9 68.6 7.5
CODA (Smith et al. 2023) 76.3 7.0 71.0 7.2
COLT (Ours) 79.8 5.8 74.7 6.4
We additionally compare the proposed COLT with popular continual learning methods on VideoToolBench test split to demonstrate the life-long tool-usage learning capability of our method.
Evaluation Metrics. To evaluate how the system retains tool-using knowledge over time, we adapt the conventional continual learning metric (Wang et al. 2024a; Chaudhry et al. 2018) to our tool-using scenario. We set up two metrics: 1) average accuracy of a specific tool denotes the overall call accuracy during the incremental learning process; 2) average forgetting is defined as the mean difference between the maximum tool call accuracy throughout the past learning process and the current tool call accuracy. A reduction in the value of average forgetting is indicative of enhanced continual learning capability. Refer to supplementary materials for the formula and detailed explanations.
Compared Methods. We set comparison experiments as follows: 1) Sequential training: training on new data with the pre-trained weights on previous tool data as initialization; 2) Rehearsal training: replay past tool data (i.e., “buffer”) and combine them with new data. Buffer size is respectively set to 10/30/50 for each tool in experiments; 3) Popular continual learning methods including L2P (Wang et al. 2022c), Dual (Wang et al. 2022b) and CODA (Smith et al. 2023).
Performance Analysis. We conduct experiments in two settings, i.e., five/ten groups with two/one tools per group. The comparison results in Table 6 reveal that our proposed COLT consistently exhibits superior performance in both average accuracy and forgetting across both settings. For example, COLT outperforms CODA (Smith et al. 2023) by 3.5% on average accuracy under the group number of five, highlighting its adeptness in mitigating the deleterious catastrophic forgetting.
Ablation Studies
We conduct extensive ablation studies to provide more insights into our proposed COLT. The experiments are conducted on MVBench (Li et al. 2023c) with the continual learning model variant COLT $_{\text{5$\times$2}}$.
Ablation on Training Strategies. Our COLT adopts a three-stage training pipeline. We conduct ablation studies on the training process by respectively skipping the second and third stages of training. The comparison results on MVBench (Li et al. 2023c) in Table 4(a) demonstrate that both stages are crucial to the final performance, e.g., skipping the second stage leads to a 1.4% drop in average scores.
Ablation on Training Loss. We use the straight-through estimator (Van Den Oord, Vinyals et al. 2017; Bengio, Léonard, and Courville 2013) for the training of stage 2 and stage 3 (c.f. Eq. (5)), which contains the quantisation loss $\mathcal{L}_{q}$ and commitment loss $\mathcal{L}_{c}$ to enables the mutual updates between the selected tool prompts and query features. The ablation results in Table 4(b) underscore the significance of both quantisation loss $\mathcal{L}_{q}$ and commitment loss $\mathcal{L}_{c}$.
Ablation on Prompt Selection. Recall that we select the most matched tool prompts based on the similarity between the textual instruction features and codebook (i.e., text-codebook). Here we ablate on the prompt selection based on the vision-codebook similarity or the average similarities of both. The results are listed in Table 4(c), which showcases that the text-codebook similarity is more reliable. That could be attributed to that it is easier for video LLMs to decide whether to invoke or which specific tool to invoke from the user instructions instead of the input visual data.
Ablation on Tool Prompt Position. We ablate on the insertion position of the tool prompt. We’ve listed three options including tool-vision-text, vision-tool-text, and vision-text-tool, demonstrating placing the tool feature before/between/after the vision and text features, respectively. The findings in Table 4(d) indicate that the performance is insensitive with regard to the tool prompt position.
Table 7: Ablations on hyper-parameters including the codebook size NN and the number of selected prompt $K$.
NN 20 30 40 50 60
AVG 48.4 50.9 51.4 51.8 51.8
$K$ 1 2 3 4 5
AVG 50.8 51.6 51.8 51.3 51.3
Figure 4: Accuracy of zero-shot video-question answering on MSVD vs. continually introduced tasks. The specific tool name for each abbreviation is available in Table 1.
Ablation on hyper-parameters. We conduct ablation studies on the codebook size NN and the number of selected prompt $K$. The comparison results are listed in Table 7. As shown, the average performance is positively correlated with the codebook size NN and reaches saturation at $T=50$. The optimal results are achieved when setting $K=3$.
Visualizations of performance vs. continually introduced tasks. To intuitively show the impact of incrementally introducing tools, we show the phased zero-shot video-question answering performance under different training strategies. Specifically, we report the accuracy score on MSVD-QA dataset in Figure 4. As shown, equipping video LLMs with dense video caption tools leads to the significant performance boost for all four training strategies. Besides, our proposed COLT shows better learning stability, i.e., less forgetfulness when faced with new tool data.
Conclusions
In this work, we present COLT which continually learns new tool-using knowledge in a data-stream scenario for general-purpose video understanding. To achieve this, we propose a learnable tool codebook where the specific tool is retrieved and activated according to the similarities with the user instructions. Due to the absence of a video-centric tool-using instruction-tuning dataset within the community, we devised VideoToolBench to address this deficiency and foster the exploration of tool-using capacities of video LLMs. Experimental results indicate that our proposed COLT adeptly invokes the necessary tools with precision, thereby achieving state-of-the-art performance on widely used benchmarks and the proposed VideoToolBench.
Appendix A1 Appendix
The outline of the appendix is as follows:
  • Dataset and Evaluation Metric Details
    • Dataset Structure of VideoToolBench
    • Prompt for Curating VideoToolBench
    • Instruction Tuning Dataset Reformat
    • Quality Control of VideoToolBench
    • Evaluation Metrics of Tool Continual Learning
  • Experimental Results
  • Broader Impact
  • Qualitative Results
Dataset and Evaluation Metric Details
Dataset Structure of VideoToolBench. Each instance in VideoToolBench contains two rounds of conversation between human and gpt. The response of gpt includes three fields of thought, action and value. One example is available in Table A1.
Table A1: Dataset structure of VideoToolBench.
An example of VideoToolBench dataset
First Round of Conversation
Human Please segment the objects in the video.
GPT <<thoughts>>
Request a process of video object segmentation based on the user’s prompt.
<<actions>>
[{‘API_\_name’: ‘video-object-segmentation’, ‘API_\_params’: {}}]
<<values>>
Sure thing! I’$l$ initiate a object segmentation model to fulfill your request. Please wait while the processing takes place.
Second Round of Conversation
Human You have successfully completed the process of video captioning and object segmentation.
GPT <<thoughts>>
I have successfully completed the process of video object segmentation. Now, it is time to update the user.
<<actions>>
<<value>>
The objects have been segmented. You can now review the results.
Prompt for Curating VideoToolBench. We collect VideoToolBench by self-instruct GPT-3.5-turbo. The prompt and in-context learning cases are shown in Table A2.
Table A2: System prompts and in-context learning examples for the generation of VideoToolBench.
Prompts for GPT-3.5-turbo to generate instruction-tuning dataset:
You are an AI assistant and you receive a set of conversations in json string format, the content of which is used as data for instruction finetune. Your task is to rephrase the dialogue to generate a new example. Note to give it in standard json format. DO NOT modify the “from” and “actions” parts. Only modify the “value” and “thoughts” part.
In-context learning examples:
{”from”: ”human”, ”value”: ”Transcribe the audio content from the provided video.”, ”actions”: []}, {”from”: ”gpt”, ”thoughts”: ”I need to find an ASR model and initiate the transcription process as per the user’s request.”, ”actions”: [{”API_\_name”: ”asr”, ”API_\_params”: {}}], ”value”: ”Certainly! I’$l$ employ an Automatic Speech Recognition (ASR) model. Please wait while the ASR model processes the video”}, {”from”: ”human”, ”value”: ”asr output: You have used an ASR model to transcribe the audio content from the provided video based on my original request.”, ”actions”: []}, {”from”: ”gpt”, ”thoughts”: ”Now that the ASR model has processed the video, I can update the user on the transcription results.”, ”actions”: [], ”value”: ”The ASR transcription you requested has been completed. Here is the transcribed text for your reference.”}
Instruction Tuning Dataset Reformat. In the second and third stage of COLT training, we reformat the instruction tuning dataset from LLaVA v1.5 (Liu et al. 2023a) and Video-ChatGPT (Maaz et al. 2023) into the thought-action-value pattern for the data format conformity. Specifically, the thought is generated by GPT indicating that the question can be answered without invoking any tools: The questions can be answered by the information in the context, without need any external tools. The action is an empty list while value is the original response.
Quality Control of VideoToolBench: To ensure the reliability and integrity of the proposed dataset, we conduct quality control from two primary aspects: data format checks and of semantic meaning checks. The format check entails verifying whether the generated data adheres to a predefined structure, specifically confirming that each entry corresponds to a two-round conversation and that the responses from the GPT model contain essential components such as thought/action/value keys. Furthermore, this phase involves confirming the consistency of tool names used throughout the dataset. Manual checks on semantic meanings are conducted to validate the contextual relevance and accuracy of the generated content, ensuring that it aligns with the intended purpose of the dataset and maintains coherence within the conversations.
Evaluation Metrics of Tool Continual Learning. We detail the metrics of average accuracy and average forgetting used in Sec. Comparisons with Continual Learning Methods.
Let $\alpha_{k,j}\in[0,1]$ denote the tool call accuracy of $j$-th tool after incrementally training on the $k$ tool data ($j\leq k$). The average accuracy of a specific tool denotes the overall call accuracy during the incremental learning process. Then the average accuracy at task $k$ is defined as follows.
$\text{AA}_{k}=\frac{1}{k}\sum_{j=1}^{k}a_{k,j}.$ (A1)
Since average accuracy does not provide any information about the forgetting profile of the continual learning process, average forgetting is introduced to bridge the gap. For a particular tool, the forgetting measure is defined as the difference between the maximum tool call accuracy throughout the past learning process and the current tool call accuracy. In particular, the forgetting for the $j$-th tool after incrementally training up to $k$ tool is as follows.
$f_{j}^{k}=\max_{l\in\{1,\cdots,k-1\}}a_{l,j}-a_{k,j},\quad\forall j<k.$ (A2)
The average forgetting of $k$-th tool is computed by normalizing against the number of tools seen previously:
$\text{AF}_{k}=\frac{1}{k-1}\sum_{j=1}^{k-1}f_{j}^{k}.$ (A3)
We report the average accuracy and average forgetting after the last tool learning.
Experimental Results
In Sec. Training, we employ an additional CLIP-initialized encoder ${f}_{\boldsymbol{\beta}}$ to extract query feature $\mathbf{H}_{\mathbf{q}}$. Then in Eq. (2), the appropriate tool prompt is selected based on the cosine similarity with $\mathbf{H}_{\mathbf{q}}$. Here we conduct the ablation study which directly uses the encoded LLM features $\mathbf{H}_{\mathbf{w}}$ for similarity computation. The comparison results in Table A3 demonstrate that the introduced CLIP feature $\mathbf{H}_{\mathbf{q}}$ leads to better performance. This might be due to the fact that LLM features are more compatible with auto-regressive language modeling, whereas CLIP features are better suited for retrieval-oriented applications.
Table A3: Ablations of the query feature source for tool prompt selection.
Query Feature Source AVG
LLM feature $\mathbf{H}_{\mathbf{w}}$ 47.4
CLIP feature $\mathbf{H}_{\mathbf{q}}$ 51.8
Broader Impact
The ability of our COLT to continually learn new tool usage without catastrophic forgetting may support the development of personalized video LLMs. Data within specific domains or pertaining to particular users can be incrementally fine-tuned, leading to an ever-evolving and personalized intelligent assistant. However, there also exists the risk that the model could be exploited for malicious purposes. Besides, it may raise privacy concerns, especially if it is deployed in surveillance systems or social media platforms.
Qualitative Results
We present additional qualitative results on MVBench (Li et al. 2023c), zero-shot video-question answering (Chen and Dolan 2011; Xu et al. 2016; Yu et al. 2019), and VideoToolBench test split in Figure A1, Figure A2, and Figure A3, respectively. Notably, our COLT precisely captures dynamic video information and generate more reasonable responses.
Figure A1: Qualitative results on MVBench.
Figure A2: Qualitative results on zero-shot video-question answering.
Figure A3: Qualitative results on the test split of VideoToolBench.